# Electrical conductivity of a random nanowire network: comparison of two-dimensional and quasi-three-dimensional models

### v1

@@ Abstract
In a recent work Tarasevich, Eserkepov, and Vodolazskaya (2026), it was shown that, within the mean-field approach, the effective electrical conductivity of such a random network of conductors is
@@
Fundamentally different dependencies of the number of contacts per conductor on the number density of conductors in the 2D and Q3D models lead within a mean-field-approach to significantly different dependencies of electrical conductivity obtained on the number density (Fig. 3). In the case where contact resistances dominate over conductor resistance, the differences in electrical conductivities obtained within the 2D and Q3D models are most significant and for the values of the number density presented in the figure reach two orders of magnitude. These differences decrease significantly with increasing parameter .
@@
In any case, a network formed by straight thin conductors randomly placed on a substrate and having uniformly distributed orientations has significantly lower connectivity than a network formed by zero-width rods. However, in the case of a two-layer network formed by conductors oriented in two mutually perpendicular directions, the differences between the 2D and Q3D models are leveled out.

### v2

@@ Abstract
In a recent work Tarasevich, Eserkepov, and Vodolazskaya (2026), it was shown that, within the mean-field approach (MFA), the effective electrical conductivity of such a random network of conductors is
@@
Fundamentally different dependencies of the number of contacts per conductor on the number density of conductors in the 2D and Q3D models lead within the MFA to significantly different dependencies of electrical conductivity obtained on the number density (Fig. 3). In the case where contact resistances dominate over conductor resistance, the differences in electrical conductivities obtained within the 2D and Q3D models are most significant and for the values of the number density presented in the figure reach two orders of magnitude. These differences decrease significantly with increasing parameter .
@@
The saturation effect of the number of contacts per conductor in a real nanowire network can be simulated using a simple two-dimensional model with memory. Consider a square domain on a plane. The domain size is . Periodic boundary conditions are used to reduce boundary effects. Segments of length are randomly and independently deposited into the domain . The segment orientations are equally probable. The segments are assigned ordinal numbers . When a new segment with number is deposited, its intersection with previously placed segments is checked. If segment intersects segment , then contact (link) between them occurs only if the condition is satisfied, where plays the role of the memory depth. Within this model, the number of contacts per conductor saturates with increasing conductor concentration (Fig. 4). When , the average number of contacts per rod tends to the value
Figure 5 compares the dependencies of the electrical conductivity on the number density of conductors obtained within the MFA and by simulation using the 2D model with memory for the case when the junction resistance dominates over the wire resistance (). The reason for the overestimated electrical conductivity within the MFA was explained in Ref. Tarasevich, Eserkepov, and Vodolazskaya, 2026.
In any case, a network formed by linear thin conductors randomly placed on a substrate and having uniformly distributed orientations has significantly lower connectivity than a network formed by zero-width rods. However, in the case of a two-layer network formed by conductors oriented in two mutually perpendicular directions, the differences between the 2D and Q3D models are leveled out.